OpenStudy (anonymous):

Derivative of 2x^(2pi)

OpenStudy (rulnick):

If derivative is wrt x, then it is 4 pi x^(2pi - 1)

geerky42 (geerky42):

\[\Large\dfrac{d}{dx} cx^n = ncx^{n-1} \\ \text{So}~\Large \dfrac{d}{dx}2x^{2\pi} = (2)(2\pi)x^{2\pi - 1} = \boxed{4\pi x^{2\pi - 1}}\]

geerky42 (geerky42):

@JS03670